How do you solve this. How much whipping cream and 2%milk should be mixed to obtain 35gallons of milk with 4% butterfat? 2%milk has 2% butterfat Whipping cream has 9% butterfat
Answer by RAY100(1637)

Let m = amt milk,,,,and c = amt of cream,,,,both in gallons
.
.02m +.09c = .04 (35)
.
but m+c =35,,,,,or m=35-c,,,,subst
.
.02(35-c) + .09 c = .04(35)
.
.7 -.02c + .09c = 1.4
.
.07c = .7
.
c= 10 gal,,,,,,,,,m=35-10=25 gal
.
.
checking,,,,,.02(25) + .09(10) = .04(35)
.
.5 + .9 = 1.4,,,,,ok
.